# Rethinking Detection Based Table Structure Recognition for Visually Rich Documents

### 01 Dec 2023 v1

@@ 2.2 Table Structure Recognition
At last, graph-based methods usually apply either detection or segmentation methods to obtain the locations of table cells and further build the relation among table cells. For instance, TGRNet xue2021tgrnet formulates the cell location detection and cell logical location prediction jointly in a multi-task architecture, which is modularized by a segmentation based method and graph convolutional network (GCN), respectively. Similarly, TSRNet li2022table proposes a unified GNN-based approach modeling table detection and table structure recognition tasks together. More specifically, TSRNet also employs a semantic segmentation module to extract primitive regions, then applies k-nearest neighbors and line-of-sight neighbors to construct the graph and further classify the graph nodes and edges to filter the noise regions, merge, and build relations. In contrast, LGPMA qiao2021lgpma proposes a Local Pyramid Mask Alignment Module and Global Pyramid Mask Alignment Module to localize table cells, which are formulated as detection and segmentation problems and can be implemented by MaskR-CNN he2017mask . To construct the structure of the table, LGPMA further proposes a pipeline of cell matching, empty cell searching, and empty cell merging using the Maximum Clique Search algorithm and rule-based methods. Besides building graph explicitly, some studies tensmeyer2019deep ; zhang2022split ; nguyen2023formerge predict the table grids or separators first, and then merge grid elements, which are also treating grid elements as graph nodes. SPLERGE tensmeyer2019deep is a typical method following this strategy consisting of a Split Model and Merge Model, in which the Split Model consists of a Row Projection Network and a Columns Projection Network to obtain the table grid, and the Merge Model is used to merge the grid cells. Similarly, SEM zhang2022split employs a segmentation model to segment columns and rows and generate the table grid with a post-processing method. After the table grid is obtained, SEM introduces an Embedder network to extract and fuse the features from textual and visual modalities. A Merger network takes the fused features from Embedder as inputs to merge the grid elements. TSRFormer-DQ-DETR WANG2023109817 leverages a DETR  carion2020end based separation line prediction model, termed DQ-DETR, to predict the reference points on separation lines, followed by a Relation Network based cell Merging module to merge grid elements.
@@ 3 Proposed Method
In this section, we first introduce the overall architecture of the proposed method, then give the details of the proposed methods, including tuning aspect ratios for the anchor boxes generation, refactoring the multi-label definition, and the proposed Spatial Attention Module.
3.1 Overall Architecture
Our proposed solution is based on Cascade R-CNN Cai2018CascadeRD , and the overall architecture is shown in Figure 4. It mainly consists of a backbone network, a region proposal network (RPN), and a series of cascade heads. The backbone network is used to extract features of the input images, which is often followed by Feature Pyramid Network (FPN) lin2017feature to extract and fuse features from different scales. The extracted features, together with the region proposals generated by the RPN, are fed into the first cascade head for the classification and regression tasks, and the regression results would be the inputs of the subsequent cascade head, as shown in Figure 4.
Since there are multiple cascade heads, all the outputs of these cascade heads are used to calculate the loss. Moreover, the final loss of the model can be defined as the sum of these Cascade Heads loss and the RPN loss, as defined by Equation 1 where $N$ is the number of cascade head. It is worth mentioning that we follow the most popular Cascade R-CNN model to use three cascade heads in our model.
$e_{cls}=PROJ(ROI\_POOL(BACKBONE(x),b))$ (2)
$e_{reg}=PROJ(ROI\_POOL(BACKBONE(x),b))$ (3)
Figure 4: Overall architecture of Cascade R-CNN.
As shown in Figure 4, there are a REG Head and a CLS Head in each cascade head for the regression and classification tasks, respectively. The input features of these REG Heads and CLS Heads $e_{cls},e_{reg}$ are extracted by applying ROI Pooling operations to the features from Backbone Network with the proposal boxes $b$, which can be defined by Equations 2 and  3. $PROJ,ROI\_POOL$, and $BACKBONE$ are the projection layer, ROI Pooling operation, and the backbone network. Therefore, for a trained model, the features of the CLS Heads $e_{cls}$ are determined by the input image $x$ and the proposal boxes $b$, meaning that a single proposal box cannot be classified into multiple classes because CLS Heads are not multi-label classifiers. As aforementioned, the TSR definition of this study is a multi-label detection problem, we refactor the problem formulation by generating pseudo-classes which will be discussed in Section 3.3. Besides, we also improve the proposal generation of the RPN network by tuning the aspect ratios, and feature generation of the backbone network by incorporating deformable convolution and Spatial Attention, which will be discussed in sections 3.2 and  3.4, respectively.
3.2 Aspect Ratio Tuning
Aspect Ratios are the parameters used in the RPN to generate anchor boxes for the two-stage detectors, which control the shape of the generated anchor boxes. Popular implementations of Cascade R-CNN, such as Detectron2 wu2019detectron2 , usually use 0.5, 1.0, and 2.0 as default values, which can work well for detecting common objects, such as the objects in COCO lin2014microsoft dataset. However, in the context of TSR, the range of aspect ratios is much larger because of the shape of the table components. For example, as shown in Figure 1(c), the aspect ratio of the last row at the table bottom is $20:1084=0.01845$, much smaller than the popular value $0.5$. Therefore, adjusting the aspect ratios to the TSR problem is important. Without proposing fancy new modules to select suitable values, we simply select the values based on the statistics of the training sets. Taking the FinTabNet dataset as an example, the aspect ratios of the defined components are shown in Figure 5. The maximum value is 140, far larger than the popular choices in common object detection. Besides, the majority of aspect ratios in Figure 5 are in the range between 1 and 60. Therefore, we extend this parameter for our proposed TSRDet as [0.0125, 0.025, 0.0625, 0.125, 0.25, 0.5, 1.0, 2.0, 4.0, 8.0, 16, 40, 80]. A further detailed parameter table is provided in Section 4. It is worth mentioning that when an aspect ratio is less than 1, its multiplicative inverse is applied to count the number of aspect ratios in Figure 5. We also didn’t further fine-tune this parameter through validation, which means that they might not be optimal. But this parameter has improved the model performance by around 2.7% as shown in Section 4.3.
Figure 5: Statistics of aspect ratio values of FinTabNet training set. When a aspect ratio is less than 1, then its multiplicative inverse is applied to count the number of aspect ratios.
3.3 Pseudo-class Generation
As mentioned earlier, this study also formulates the TSR problem as detecting six components, namely Table, Column, Row, Spanning Cell, Project Row Header and Column Header, as shown in Figure 1(c) and Figure 3. In the study in smock2022pubtables , all these six components are treated independently. Formally, the ground truth $\mathbf{Y}$ of this definition for each image is a set of tuples containing bounding boxes and their corresponding labels, as defined by Equation 5 where $b_{i},c_{i}$ are the $i$th bounding box and class, and values from 00 to $5$ are the Table, Column, Row, Spanning Cell, Project Row Header and Column Header, respectively. This definition is a multi-label problem for the classification task of a detection model, because some table components, such as Column Header and Row can share identical bounding boxes, as shown in Figure 3. This multi-label definition is challenging for typical two-stage detectors, because the classifiers in these detectors are designed for single-label classification.
@@ 3.3 Pseudo-class Generation
Therefore, in this study, considering the observation that the defined Project Row Headers are all Rows at the same time, we only keep the Project Row Headers samples during the training. Since some Column Headers can share identical bounding boxes with corresponding Rows, we derive a pseudo class for these overlapped samples and remove the original overlapped samples. Therefore, during the training stage, we refactor the ground truth for each image to the regular single-label classification, as defined by Equation 4 where values 00 to $6$ are the Table, Column, Row, Spanning Cell, Project Row Header, Column Header and the Pseudo Class, respectively. During the testing stage, the results of Project Column Header are duplicated once to generate their corresponding prediction Rows, and the results of the pseudo-class are duplicated twice to generate the corresponding prediction Rows and Headers, so that we can still follow the formulation defined by Equation 4 to evaluate the model performance.

### 10 Jan 2024 v2

@@ 2.2 Table Structure Recognition
At last, graph-based methods usually apply either detection or segmentation methods to obtain the locations of table cells and further build the relation among table cells. For instance, TGRNet xue2021tgrnet formulates the cell location detection and cell logical location prediction jointly in a multi-task architecture, which is modularized by a segmentation based method and graph convolutional network (GCN), respectively. Similarly, TSRNet li2022table proposes a unified GNN-based approach modeling table detection and table structure recognition tasks together. More specifically, TSRNet also employs a semantic segmentation module to extract primitive regions, then applies k-nearest neighbors and line-of-sight neighbors to construct the graph and further classify the graph nodes and edges to filter the noise regions, merge, and build relations. In contrast, LGPMA qiao2021lgpma proposes a Local Pyramid Mask Alignment Module and Global Pyramid Mask Alignment Module to localize table cells, which are formulated as detection and segmentation problems and can be implemented by MaskR-CNN he2017mask . To construct the structure of the table, LGPMA further proposes a pipeline of cell matching, empty cell searching, and empty cell merging using the Maximum Clique Search algorithm and rule-based methods. Besides building graph explicitly, some studies tensmeyer2019deep ; zhang2022split ; nguyen2023formerge predict the table grids or separators first, and then merge grid elements, which are also treating grid elements as graph nodes. SPLERGE tensmeyer2019deep is a typical method following this strategy consisting of a Split Model and Merge Model, in which the Split Model consists of a Row Projection Network and a Columns Projection Network to obtain the table grid, and the Merge Model is used to merge the grid cells. Similarly, SEM zhang2022split employs a segmentation model to segment columns and rows and generate the table grid with a post-processing method. After the table grid is obtained, SEM introduces an Embedder network to extract and fuse the features from textual and visual modalities. A Merger network takes the fused features from Embedder as inputs to merge the grid elements. TSRFormer-DQ-DETR WANG2023109817 leverages a DETR  carion2020end based separation line prediction model, termed DQ-DETR, to predict the reference points on separation lines, followed by a Relation Network based cell Merging module to merge grid elements. Since these graph-based models identify the graph nodes first, defining a cell-type classification task is necessary if they want to provide information regarding header cells.
3 Rethinking Detection-based TSR models
3.1 Preliminaries
Since most existing detection-based TSR models are based on two-stage and transformer-based detectors, we use Cascade R-CNN Cai2018CascadeRD and Sparse R-CNN sun2021sparse as two examples of these two types of detectors and briefly review their critical designs in this section.
3.1.1 Cascade R-CNN
Cascade R-CNN Cai2018CascadeRD is a typical two-stage detection model containing a Backbone Network, a Region Proposal Network (RPN), and a series of Cascade Heads, as shown in Figure 1. The RPN is the first regression step of a two-stage detection model responsible for generating region proposals. More specifically, a set of predefined anchor boxes are defined and slides across the feature map to generate the fix-length of feature vectors for the classification and regression tasks in the RPN ren2015faster . The classification task classifies anchor boxes into object and background, and the regression task coarsely regresses the anchor boxes to generate higher-quality region proposals. Since the RPN only coarsely classifies and regresses the anchor boxes, the parameters of defining anchor boxes play a key role in the performance of the RPN, such as the number of anchor boxes, the aspect ratios of anchor boxes, and the scales of applied feature maps.
Figure 1: Overall architecture of Cascade R-CNN.
The Backbone Network is used to extract features of the input images, which is often followed by Feature Pyramid Network (FPN) lin2017feature to extract and fuse features from different scales. The extracted features, together with the region proposals generated by the RPN, are fed into the first cascade head for the classification and regression tasks, and the regression results would be the inputs of the subsequent Cascade Head, as shown in Figure 1.
Since there are multiple Cascade Heads, all the outputs of these Cascade Heads are used to calculate the loss. Moreover, the final loss of the model can be defined as the sum of these Cascade Heads loss and the RPN loss, as defined by Equation 1, where $N$ is the number of Cascade Heads. It is worth mentioning that we follow the most popular Cascade R-CNN model to show three Cascade Heads in Figure 1. Each Cascade Head has a REG Head and a CLS Head for the regression and classification tasks, respectively. The input features of these REG Heads and CLS Heads $e_{cls},e_{reg}$ are extracted by applying ROI Pooling operations to the features from Backbone Network with the proposal boxes $b$, which can be defined by Equations 2 and  3 where $PROJ,ROI\_POOL$, and $BACKBONE$ are the Projection layer, ROI Pooling operations, and the Backbone Network. Therefore, for a trained model, the input features of the CLS Heads $e_{cls}$ are determined by the input image $x$ and the proposal boxes $b$, meaning that a single proposal box cannot be classified into multiple classes because CLS Heads are not multi-label classifiers.
$e_{cls}=PROJ_{cls}(ROI\_POOL(BACKBONE(x),b))$ (2)
$e_{reg}=PROJ_{reg}(ROI\_POOL(BACKBONE(x),b))$ (3)
3.1.2 Sparse R-CNN
Figure 2: Overall architecture of Sparse R-CNN.
Sparse R-CNN is a popular end-to-end transformer-based detection model. Similar to Cascade R-CNN, Sparse R-CNN also employs a cascade architecture containing a series of Dynamic Heads, as shown in Figure 2. In each Dynamic Head, an ROI Pooling layer is applied to extract features from the feature map based on the given proposal boxes, and the extracted features, together with the learnable proposal features, are fed to the Dynamic Instance Interactive Head to generate final features for the classification and regression tasks. Therefore, the features fed into CLS Head and REG Head of each Dynamic Head can be defined as Equations 4 and  5, where $BACKBONE$, $DYN\_HEAD$, and $PROJ$ are the Backbone Network, Dynamic Instance Interactive Head and the Projection layer, respectively, and $x$, $b$, $f$ are the input image, the proposal boxes and the learnable proposal features. It is worth mentioning that Sparse R-CNN does not use any RPN network to generate regional proposals. Instead, it proposes to use a set of learnable proposal boxes paired with a set of learnable features, in which learnable proposal boxes can be initialized by some pre-defined methods, such as image size initialization, random initialization, and grid initialization. Once the model is trained, the proposal boxes can be treated as an identical value, such as the box of image size, and their classification and regression results are mainly determined by their corresponding learnable proposal features $f$ and the input image $x$. Therefore, for a multi-label detection problem, when objects belonging to different classes can share an identical box, the learnable proposal features can be different for these objects, making it possible for Sparse R-CNN to deal with multi-label detection tasks.
$e_{cls}=PROJ_{cls}(DYN\_HEAD(ROI\_POOL(BACKBONE(x),b),f))$ (4)
$e_{reg}=PROJ_{reg}(DYN\_HEAD(ROI\_POOL(BACKBONE(x),b),f))$ (5)
3.2 Rethinking Problem Formulations
As aforementioned in section 1, there have been many detection-based solutions siddiqui2019deeptabstr ; hashmi2021guided ; fernandes2023tablestrrec ; xiao2022efficient ; smock2022pubtables with different problem formulations that either oversimplified the TSR task or ignored its multi-label characteristic. More specifically, following image-to-sequence TSR models, a detection-based TSR model should be able to fully reconstruct the structure of both regular and spanning table cells, as well as provide information regarding header cells. However, studies siddiqui2019deeptabstr ; hashmi2021guided formulate the problem as only detecting columns and rows, making them impossible to deal with spanning cells and identify header cells. TableStrRec fernandes2023tablestrrec further extend the formulation by defining regular column, regular row, irregular column, and irregular row so that the spanning cell can be inferred from these four types of components, as shown in Figure 2(a). But still, this formulation cannot provide information regarding header cells, which is still over-simplied the TSR task. Study xiao2022efficient simplifies the formulation of in PubTables1M smock2022pubtables , directly detecting table, column, row, and spanning cell, as shown in Figure 2(b), which is another formulation ignoring header cells’ information. Besides, these three formulations treat the Projected Row Header as a regular row, resulting in over-simplified table structures. By contrast, PubTables1M smock2022pubtables defines six types of components, including Table, Column, Row, Spanning Cell, Column Header, and Projected Row Header, as shown in Figures 2(c) and  2(d), which can provide as much structure information as image-to-sequence TSR models. However, this formulation does not consider that some Column Headers and Projected Row Headers can share identical bounding boxes with corresponding Rows. For example, as shown in Figure 2(c), the bounding boxes of the two Projected Row Headers can also be classified as Rows. Similarly, as shown in Figure 2(d), the Column Header’s bounding box is also the Row’s bounding box. Therefore, the problem definition of study smock2022pubtables is a multi-label detection problem, which is must be considered when we choose and design detection models. It is worth mentioning that all these problem formulations use extracted table images as inputs. Even though many studies cascadetabnet2020 ; xiao2023revisiting ; xiao2023table have achieved very promising performance on the Table Detection (TD) task, it is still difficult to guarantee that all the table content can be fully included in the detection results. Therefore, in practice, the detected bounding boxes of tables from the TD model are often padded with extra pixels, making it necessary to define a Table component for TSR.
(a) Four types of defined table components in TableStrRec fernandes2023tablestrrec . This definition can infer Spanning Cells but cannot provide Column Header information. Besides, the Projected Row Headers are treated as regular Rows, which can lead to a wrong structure.
(b) Four types of defined table components in the study xiao2022efficient . This definition cannot provide Column Header information, and the Projected Row Headers are treated as regular Rows, which can lead to a wrong structure.
(c) Six types of defined table components in PubTables1M smock2022pubtables . The defined Projected Row Headers in this sample share identical bounding boxes with two corresponding rows.
(d) Six types of defined table components in PubTables1M smock2022pubtables . In this sample, the defined Column Header shares an identical bounding box with a defined Row. It is worth mentioning that there are no Spanning Cell and Projected Row Header in this sample.
3.3 Revisiting Region Proposal Generation
As aforementioned in section 3.1, the parameters of generating anchor boxes in the RPN play a key role in two-stage detection models, while transformer-based detection models, such as DETR and Sparse R-CNN, use learnable queries or proposals without the need to tune the RPN. If we choose two-stage detection models, such as Cascade R-CNN which is used in TableStrRec fernandes2023tablestrrec , we need to identify the difference between the TSR detection problem and widely discussed common object detection problem, because the default settings of detection frameworks, such as Detectron2 wu2019detectron2 and MMDetection mmdetection are often tuned on COCO lin2014microsoft dataset. Therefore, we compare the statistics of the COCO dataset with a popular TSR dataset, FinTabNet zheng2020global , regarding the number of objects in each image and the aspect ratios of objects. More specifically, the COCO training set contains 118287 images and 860001 target objects, resulting in an average of 7.27 objects in each image, while the FinTabNet training set contains 78537 images, 1628298 target objects, resulting in an average of 20.73 objects in each image. Besides, the aspect ratios of objects in these two dataset are also very different, as shown in Figure 4. The vast majority of target objects in the COCO training set have aspect ratios between 1 and 10, while objects in the FinTabNet training set have much larger aspect ratios. Therefore, we need to consider these differences when tuning the parameters of RPN if we apply a two-stage object detection model for the TSR task, such as increasing the number of region proposals and adjusting the aspect ratios of anchor boxes. On the other hand, transformer-based detection models, such as Sparse R-CNN and DETR, can alleviate the issues caused by these differences intrinsically because they use learnable queries (learnable proposals) instead of an RPN, as discussed in section 3.1. However, increasing the number of learnable queries for each image might be also useful for transformer-based detection models because TSR datasets contain more objects than common object detection datasets.
(a) Aspect ratios of the FinTabNet training set.
(b) Aspect ratios of the COCO training set.
Figure 4: Statistics of aspect ratio values of COCO and FinTabNet training sets. When an aspect ratio is less than 1, its multiplicative inverse counts the number of aspect ratios.
3.4 Rethinking Detection and TSR Metrics
As mentioned in sections 1 and  2, detection-based TSR models need a deterministic rule-based post-processing method to transform the detected objects into structured sequences. Existing studies siddiqui2019deeptabstr ; hashmi2021guided ; fernandes2023tablestrrec ; xiao2022efficient usually use the detection performance to evaluate the model performance before applying the post-processing method. However, the detection metrics are not aligned with cell-level TSR metrics. We use COCO lin2014microsoft and TEDS zhong2020image metrics as examples for further analysis in this section. The COCO metrics employ mean Average Precision (mAP) to evaluate the model performance, which can be defined by Equation 6 where $N$, $precision_{i}(r)$ and $dr$ in Equation 6 are the number of classes, and the precision at a given recall level $r$ for class $i$. In practice, the precision-recall curves in COCO metrics are computed for each class at a series of IoU thresholds, and the integral of $precision_{i}(r)$ often is approximated by the discrete sum. The IoU score can be defined by Equation  7, where $A\cap B$, $A\cup B$ are the intersection and union of bounding boxes $A$ and $B$. In practice, in many studies, mAP is represented by AP and calculated by averaging the mean precision scores of all categories at IoU thresholds from 0.5 to 0.95 with 0.05 intervals. AP50, AP75 are the mean precision scores of all categorizes at IoU thresholds 0.5 and 0.75, respectively. Therefore, COCO metrics are IoU-based evaluation metrics. By contrast, TEDS can be defined by Equation 8, where EditDist is the tree-edit distance, and $T$ is the number of nodes in the tree, meaning that TEDS is not correlated with IoU scores.
[…]
$IoU=\frac{A\cap B}{A\cup B}$ (7)
Figure 5: A sample from the FinTabNet dataset with ground truth boxes larger than the minimum bounding boxes for table structure. We only show the annotations of Columns for simplicity.
On the other hand, TSR datasets usually use a canonicalization procedure smock2022pubtables or annotate the bounding boxes following the lines in tables, which makes the ground truth boxes larger than the minimum box that can recover the structure of the table. Figure 5 shows an example from the FinTabNet dataset, whose ground truth boxes are larger than the minimum bounding boxes for table structure. Considering the four prediction boxes in Figure 5, since the prediction 1 is smaller than the minimum box for table structure, and the prediction2 can cover all content of the minimum box for table structure and has a larger IoU with the ground truth box, prediction 2 can lead to better performance regarding both COCO and TEDS metrics than prediction 1. By contrast, prediction 3 has a larger IoU with the ground truth box than prediction 2, which can lead to better detection performance. However, when it comes to TEDS, prediction 3 cannot show any superiority compared to prediction 2, because both of them can cover the minimum box for table structure. When we compare prediction 2 and prediction 4, prediction 4 has a larger IoU with the ground truth box, making it better on detection performance, but it loses information of the row, making its performance in TEDS worse than prediction 2. Therefore, because of the definitions of COCO and TEDS metrics and the procedure of annotating datasets, a detection-based TSR model might be over-optimized towards detection performance without increasing the TEDS performance and sometimes can decrease the TEDS performance.
3.5 Rethinking Feature Extraction
As mentioned in section 1, deformable convolution dai2017deformable has been applied in detection-based TSR fernandes2023tablestrrec ; siddiqui2019deeptabstr and other related solutions siddiqui2018decnt ; mondal2023dataset , demonstrating its effectiveness in improving detection performance. Deformable convolution uses a learnable grid offset to sample the grid points from the feature map and then apply the convolution operation to the sampled grid points, as defined by Equation 9,
$\mathbf{z}_{p_{0}}=\sum_{p_{n}\in R}{w(p_{n})x(p_{0}+p_{n}+\Delta p_{n})}$ (9)
where $p_{0}$ is the location on the output feature map $\mathbf{z}$, $p_{n}$ is the $n$th grid point in grid $R$, and $\Delta p_{n}$ is the $n$th learnable offset. Since the offset $\Delta p_{n}$ applied to the deformable convolution is usually obtained by a regular convolution with small kernels, such as a 3 * 3 kernel, it can only improve the local feature instead of building long-range dependencies. However, building the long-range dependencies for the TSR task is important because of the characteristics of table components. More specifically, different parts of a single table component are often sparsely distributed across the table instead of a single area of compact pixels like common objects. Figure 6 shows a sample with its Row annotations. Taking the first Row as an example, as shown in Figure 6, it mainly contains three parts, which are distributed sparsely, and there is a large space between the first part and the second part, even they all belong to a single target component. Therefore, it is important to build long-range dependencies together with improving local features, such as applying deformable convolution. And over-optimized local features, such as merely applying deformable convolution might degrade the performance regarding the TEDS.
Figure 6: A sample from the FinTabNet dataset. We only show its Row annotations for simplicity. The first Row in this Figure contains three major parts numbered 1 to 3.
@@ 4 Proposed Method
In this section, we demonstrate how to fill the performance gap between detection-based and other types of TSR models by applying very simple methods to tailor the Cascade R-CNN model based on our analysis and findings in the previous sections. We first introduce the our problem formulation, then give the details of the proposed methods, including adjusting the parameters of the RPN, applying deformable convolution and introducing Spatial Attention Module.
4.1 Proposed Problem Formulation
As mentioned in sections 3.1 and  3.2, the definition of PubTables1M smock2022pubtables can provide as much information as other types of solutions and is a multi-label detection problem, which is challenging for two-stage detectors. Therefore, we follow PubTables1M to define six table components: Table, Column, Row, Spanning Cell, Projected Row Header, and Column Header, and transform the formulation into a single-class detection problem. More specifically, we remove the Rows that share their bounding boxes with the Projected Row Header, as shown in Figure 6(a), and use a Pseudo Class to replace the Rows and Column Headers when they share identical bounding boxes, as shown in Figure 6(b). It is worth mentioning that only the Row, Projected Row Header, and Column Header are shown because the Table, Column, and Spanning Cell are the same as PubTables1M. These two samples are also in Figures 2(c) and  2(d), which show their original definition in PubTables1M.
Formally, the ground truth $Y$ in PubTable1M’s definition for each image is a set of tuples containing bounding boxes and their corresponding labels, as defined by Equation 10, where $b_{i},c_{i}$ are the $i$th bounding box and its class, and values from 00 to $5$ are the defined Table, Column, Row, Spanning Cell, Projected Row Header and Column Header, respectively.
@@ 4.1 Proposed Problem Formulation
By contrast, in this study, considering the observation that the defined Projected Row Headers are all Rows at the same time, we only keep the Projected Row Headers samples during the training. Since some Column Headers can share identical bounding boxes with corresponding Rows, we derive a pseudo class for these overlapped samples and remove the original overlapped samples. Therefore, during the training stage, we refactor the ground truth for each image to the regular single-label classification, as defined by Equation 11, where values 00 to $6$ are the Table, Column, Row, Spanning Cell, Projected Row Header, Column Header and the Pseudo Class, respectively. During the testing stage, the results of Project Column Header are duplicated once to generate their corresponding prediction Rows, and the results of the pseudo-class are duplicated twice to generate the corresponding prediction Rows and Headers, so that we can still follow the formulation defined by Equation 10 to evaluate the model performance.
@@
(a) An example of our problem formulation. In this example, we remove two Rows because their bounding boxes are identical with two Projected Row Headers.
(b) An example of our problem formulation. In this example, we derive a Pseudo Class because its bounding box simultaneously belongs to a Row and a Column Header.
Figure 7: Examples of our proposed problem formulation. Since the definitions of Table, Column, and Spanning Cells are same with PubTables1M, only Row, Column Header and Projected Row Header are showed for simplicity.
4.2 Tuning Parameters of RPN
As mentioned in sections 3.1 and 3.3, regional proposal generation is a critical step in two-stage detectors, which need to be carefully considered for the TSR problem. Therefore, we adjust Aspect Ratios and increase the number of generated regional proposals for our tailored model. More specifically, aspect ratios control the shape of the generated anchor boxes. Popular implementations of Cascade R-CNN, such as Detectron2 wu2019detectron2 , usually use 0.5, 1.0, and 2.0 as default values, which can work well for detecting common objects, such as the objects in COCO lin2014microsoft dataset. However, in the context of TSR, the range of aspect ratios is much larger because of the shape of the table components, as discussed in section 3.3. Without proposing fancy new modules to select suitable values, we simply select the values based on the statistics of the training sets. Taking the FinTabNet dataset as an example, the aspect ratios of the defined components are shown in Figure 4. The maximum value is 140, far larger than the popular choices in common object detection. Besides, the majority of aspect ratios in Figure 4 are in the range between 1 and 60. Therefore, we extend this parameter for our proposed model as [0.0125, 0.025, 0.0625, 0.125, 0.25, 0.5, 1.0, 2.0, 4.0, 8.0, 16, 40, 80]. A further detailed parameter table is provided in section 5. It is worth mentioning that when an aspect ratio is less than 1, its multiplicative inverse is applied to count the number of aspect ratios in Figure 4. We also didn’t further fine-tune this parameter through validation, which means that they might not be optimal. But this parameter has improved the model performance by around 2.7% as shown in section 5.3. Besides, since increasing the number of proposals has been applied in existing studies fernandes2023tablestrrec , and demonstrated its effectiveness, we increase it for both the base Cascade R-CNN and our proposed model.